OpenStudy (anonymous):

The facing pages of a book are consecutive integers. The product of these page numbers is 306. Find the page numbers

OpenStudy (cathyangs):

17 and 18. their product is 306.

OpenStudy (anonymous):

Let the numbers be x, x + 1. x(x + 1) = 306 x^2 + x = 306 x^2 + x - 306 = 0

OpenStudy (cathyangs):

I took the square root of 306, which is 17.4928... and the floor and ceiling integers are the answers.

OpenStudy (anonymous):

You should solve it using the formula for quadratic equation and like cathyangs says, you will get x = 17 and x + 1 = 18

OpenStudy (anonymous):

The formula for quadratic equation is